1.     The owner of a theme park would like to know the average amount of time visitors spend there. To do this, he asked 25 of them and found out that they stayed for an average of 123.5 minutes with a standard deviation of 10.5 minutes. With a confidence level of 95 percent, determine the confidence interval estimate of the population mean.


1
Expert's answer
2022-05-31T13:04:58-0400

Confidence interval:

"\=x\pm z\frac{s}{\sqrt{n}}"

z is the chosen Z-value (1.96 for 95% from t-table)

"123.5-(1.96)(\frac{10.5}{\sqrt{25}})<\mu<123.5+(1.96)(\frac{10.5}{\sqrt{25}})"

"119.384<\mu<127.616"

95% confidence interval is (119.384, 127.616)